if 3 times dan's age is decreased by 10 more than twice his age, the result is 22. what is his age?
Answer by checkley77(12844)

3x-(10+2x)=22
3x-10-2x)=22
x=22+10
x=32 ans. foe Dan's age.
Proof:
3*32-(10+2*32)=22
96-(10+64)=22
96-74=22
22=22